OpenStudy (anonymous):

The rectangular top of a small box has dimensions of 140 mm by 230 mm. How long (in centimeters) would a ribbon need to be to go all the way around the edge of the top of the box? 740 cm 37 cm 74 cm 7,400 cm

OpenStudy (anonymous):

740 cm

OpenStudy (nurali):

This will be 2L + 2W. Hence, it is 2(140) + 2(230)= 280+460=740mm. Next, convert this into cm. Remember that 1cm = 10 mm. So to convert mm to cm, you divide by 10. 740/10 = 74 cm.

OpenStudy (anonymous):

oops forgot the conversion... thanks @Nurali

OpenStudy (anonymous):

Thanks

OpenStudy (nurali):

My Pleasure.